OpenStudy (anonymous):

Can someone help me with finding the points on the graph at which the tangent line has a slope of 1. y=1/3x^3-2x^2+4x+1

OpenStudy (anonymous):

I know I have to take the derivative of this. Which i believe will be Y'=x^2-4x+4

OpenStudy (amistre64):

correct, and that formula there defines the slope of the tangent line at any given value of x what would you spose we do with it if we want to know which points have a tangent line slope of 1?

OpenStudy (anonymous):

set it equal with a slope of 1?

OpenStudy (amistre64):

yep, sooo 1=x^2-4x+4 is a quadratic equation from algebra ... can you find the values of x for this?

OpenStudy (amistre64):

or if you notice the complete squareness of the right side ... 1 = (x-2)^2 ; when x = +-1 + 2 , correct?

OpenStudy (anonymous):

yes sorry I'm a little slow I have to right everything out. but yes i got the same thing (x-2)(x-2)=1 and then splitting it up i got x=2 but i didn't get x=-1

OpenStudy (amistre64):

when x=2, y' = 0 in order to adjust the value of x, such that y' = 1 .... x = 2-1, or x = 2+1

OpenStudy (anonymous):

oh okay i see

OpenStudy (amistre64):

now that we know for what values of x we get a slope of "1" ... x = 1 or x=3 lets find our y values for the points they want. y = f(x); determine the value of f(1) and f(3) to define the points: (1,f(1)) and (3,f(3))

OpenStudy (anonymous):

so pluging in the x values we get y'=1^2+4(1)+4 which makes y=1and y'=3^2-4(3)+4 which would also make y= 1 but i don't think that's right

OpenStudy (anonymous):

or do i put it into the original equation? and not y'?

OpenStudy (amistre64):

the original equation defines points the derivative of the equation defines slopes so we determined the values of x to define a slope of 1, now we would want to use the original to define the points where the slope is 1

OpenStudy (anonymous):

oh okay so its going to be (1,10/3) for the girst point and for the second you get (3,4)

OpenStudy (amistre64):

ill trust you on those ;)

OpenStudy (anonymous):

i ment first ha sorry.

OpenStudy (anonymous):

okay thanks so much i really appriciate it

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

thanks so much